18. Solve by completing the square: a. x² - 10x = -8 b. 3x² +6x = 72
Answer #1

а - 10 - - 8 х -Jox + 8 = 0 (a+b) = at 2.4. А a-b)? ) = ? -2 -a A+32 2 = 0 - х*- 2•x.s +s-s*+ 8 =0 » (x-s) - 2S +8 = 0 = (x-s